A particles of mass m = 5 units is moving with a uniform speed v = 3 * root 2 present in the XOY plane along the line Y = X + 4. What is the magnitude of the angular momentum of the particle about the origin?
1
Expert's answer
2016-01-27T08:54:54-0500
Angular momentum of particle L=mv*r_⊥
m=5units
v=3√2
r_⊥- length of the altitude from origin to (y=4+x)
r_⊥=2√2, then
L=mv*r_⊥=5*3√2*2√2=60
answer: 60